OpenStudy (anonymous):

The admission fee at a small fair is $1.50 for children and $4.00 for adults. On a certain day, 2200 people entered the fair and $5050 was collected. How many children and how many adults went to the fair?

OpenStudy (anonymous):

x+y=2200 , 1.5 x + 4 y = 5050 ===> 1.5x+4(2200-x)=5050 ===> x= 1500 ( children) ==> adults = 2200 -1500=700